OpenStudy (anonymous):

Each set of four ordered pairs below defines vertices, in counterclockwise order, of a quadrilateral. Determine if the quadrilateral is a parallelogram.

OpenStudy (anonymous):

{(-2,-3), (-3,-6), (1,-2), (2,1)}

OpenStudy (radar):

Since slope m is equal to (y1-y2)/(x1-x2) lets check this out. (-3-(-6))/(-2-(-3)) =3/1=3 (-6-(-2))/(-3-1)=-4/-4=1 (-2-1)/(1-2)=-3/-1=3 If the next one is 1 then it is a parallelogram. (1-(-3)/2-(-2))=4/4=1

OpenStudy (anonymous):

It says all i should put for the answer is Yes or No... So i am guessing the answer is yes?

OpenStudy (radar):

Yes the opposite sides have same slope so are parallel

OpenStudy (anonymous):

Ok thanks a lot!

OpenStudy (radar):

You're welcome.